The acceleration of Point A is defined by the relation a = -1.8sinkt , where a and t are expressed in m/s2 and seconds, respectively, and k = 3 rad/s. Knowing that x = 0 and v = 0.6 m/s when t = 0, determine the velocity and position of Point A when t = 0.5 s.
1
Expert's answer
2020-04-26T18:50:55-0400

Given:

"a=-1.8\sin(3t){\:\rm m/s^2},\\
x(0)=0{\:\rm m},\\
v(0)=0.6{\:\rm m/s},\\
t=0.5\:\rm s."

The velocity of a point


"v(t)=\int a(t)dt=\\
=-1.8\int \sin(3t)dt=0.6\cos(3t)+C""v(0)=0.6\cos(0)+C=0.6+C=0.6,\; C=0."

Hence,

"v(t)=0.6\cos(3t),\\
v(0.5)=0.6\cos(3\times 0.5)=0.04\:\rm m/s."

The position of a point

"x(t)=\int v(t)dt=\\
=0.6\int \cos(3t)dt=0.2\sin(3t)+C""x(0)=0.2\sin(0)+C=C=0,\; C=0."

Hence,

"x(t)=0.2\sin(3t),\\
x(0.5)=0.2\sin(3\times 0.5)=0.2\:\rm m."